# Study of the decay pattern of $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state

### v1

@@ ACKNOWLEDGMENTS
The authors appreciate the useful discussion with Feng-Kun Guo, Jia-Jun Wu, Xiong-Hui Cao, Shu-Ming Wu, Zhen-Hua Zhang, Zhao-Sai Jia, and Xu Zhang.
@@ Appendix A Isospin conventions
Let $I$ being an isospin operator in the Hilbert space, and its corresponding isospin transform matrix is $\bm{\tau}/2$. As discussed in the Ref. [77], we know that due to the fact that $\hat{C}$ does not commute with the step operators $I_{1}\pm iI_{2}$, the relative phases of the antiparticle states will not satisfy the following requirement
@@
Note that the vertices of $V\kappa\bar{\kappa}$ do not appear as a physical process, hence it could not be directly determined from experiments. We know that the structure of the scalar is still a subject of great debate, it might be a conventional quark model state, a molecular state composed of two pseudoscalars or a four quark state, or even a mixture of these.
If $\kappa$ is more likely to be a molecular state consisting of $K\pi$, its couplings to other state might be described as follows: the $\kappa\to V\kappa/S\kappa$ interaction can be viewed as one of the pseudoscalar components inside $\kappa$ interacting with the vector or scalar, while the other acts approximately as a spectator. In this point of view, for the example of $V=\rho$, the $\kappa\kappa\rho$ vertex could be approximated to a kind of combination of the $\rho\pi\pi$ and $\rho K\bar{K}$ couplings by guaranteeing that the $K\pi$ is projected to the isospin state of $\kappa$. In the molecular picture, the flavor wave function of $\kappa$ can be expressed as
@@
The magnitude of the coupling $g_{\phi K\bar{K}}$ could be extracted by the decay of $\phi\to K\bar{K}$ with the fact that the $K\bar{K}$ channel almost saturate the width of $\phi$. Based on $g_{\phi K\bar{K}}$ we could obtain the $g_{\rho K\bar{K}}$ and $g_{\omega K\bar{K}}$ coupling by the flavor SU(3) symmetry. By the Lagrangian with SU(3) symmetry, the vertex coupling constants are connected as following [62]:
$g_{VPP}=g_{\phi K^{+}K^{-}}$ $=$ $g_{\phi K^{0}\bar{K}^{0}}=4.47e^{i\theta_{VPP}},$ (36)
@@
If regard $\sigma$ as a $\pi\pi$ molecular state, its wave function can be written as
@@
In the same way, one of the components inside the $\sigma$ is assumed to act as a spectator in the interaction, then the coupling of $\sigma\kappa\bar{K}^{*}$ vertex can be approximated to the coupling of $K^{*}K\pi$ vertex. For example, considering the respective wave functions of $\sigma$ and $\kappa$ the coupling $g_{\sigma\kappa^{+}K^{*-}}$ can be expressed as
@@ Appendix C Issues about the $\eta$-$\eta^{\prime}$ mixing
Incorporating the mixing, a U(3) flavor symmetry matrix which is the sum of the octet and singlet matrices could be adopted as shown in [78], i.e.,

### v2

@@ ACKNOWLEDGMENTS
The authors appreciate the useful discussions with Feng-Kun Guo, Jia-Jun Wu, Xiong-Hui Cao, Shu-Ming Wu, Zhen-Hua Zhang, Zhao-Sai Jia, and Xu Zhang.
@@ Appendix A Isospin conventions
Let $I$ being an isospin operator in the Hilbert space, and its corresponding isospin transform matrix is $\bm{\tau}/2$. As discussed in Ref. [77], due to the fact that $\hat{C}$ does not commute with the step operators $I_{1}\pm iI_{2}$, the relative phases of the antiparticle states will not satisfy the following requirement
@@
We assume $\kappa$ to be a molecular state consisting of $K\pi$, its couplings to other state can be described as follows: the $\kappa\to V\kappa/S\kappa$ interaction can be viewed as one of the pseudoscalar components inside $\kappa$ interacting with the vector or scalar, while the other component acts approximately as a spectator. For example, the $\kappa\kappa\rho$ vertex can be approximately related to the $\rho\pi\pi$ and $\rho K\bar{K}$ couplings by guaranteeing that the $K\pi$ is projected to the isospin state of $\kappa$. In the molecular picture, the flavor wave function of $\kappa$ can be expressed as
@@
Based on the $g_{K^{*0}K^{+}\pi^{-}}$ we can obtain the $g_{\rho K\bar{K}}$ and $g_{\omega K\bar{K}}$ couplings by the SU(3) flavor symmetry. By the Lagrangian with SU(3) symmetry, the vertex coupling constants are connected as following [62]:
$g_{VPP}$ $=$ $g_{K^{*0}K^{+}\pi^{-}}$ (36)
@@
Regarding the $\sigma$ as a $\pi\pi$ molecular state, its wave function can be written as
@@
Similarly, one of the components inside the $\sigma$ is assumed to act as a spectator in the interaction, then the coupling of $\sigma\kappa\bar{K}^{*}$ vertex can be approximately related to the coupling of $K^{*}K\pi$ vertex. For example, based on the wave functions of $\sigma$ and $\kappa$, the coupling $g_{\sigma\kappa^{+}K^{*-}}$ can be expressed as
@@ Appendix C Issues about the $\eta$-$\eta^{\prime}$ mixing
Incorporating the mixing, a U(3) flavor symmetry matrix which is the sum of the octet and singlet matrices can be adopted as shown in [78], i.e.,